An unknown amount of a compound with a molecular mass of 274.11 g/mol 274.11 g/mol is dissolved in a 10 mL volumetric flask. A 1.00 mL aliquot of this solution is transferred to a 25 mL volumetric flask, and enough water is added to dilute to the mark. The absorbance of this diluted solution at 325 nm 325 nm is 0.523 0.523 in a 1.000 cm cuvet. The molar absorptivity for this compound at 325 nm 325 nm is ? 325 =6421 M −1 cm −1 . ϵ325=6421 M−1cm−1. What is the concentration of the compound in the cuvet?

Chemistry
Answer #1

The Molar Mass of the compound= 274.11g/mol, absorbance =0.523, Molar absorptivity€325 = 64211-1cm-1

Using Beer-Lambert's law, we get -

A= EC/,

where, A= Absorbance, \epsilon- Molar Absorptivity, c-concentration of the solution in the cuvette, l- length of the cuvette=1 cm.

..C= A/1 = 0.523/(6421M .cm-1 x 1cm) = 8.14 x 10-5M.